OpenStudy (anonymous):

you have 100.00 to buy some fireworks at the local store. it cost 30.00 for a set of cherry bomba and 25.00 for a stack of roman candles. let x represent the number of cherry bombs you buy. let y represent the number of roman candley you buy

OpenStudy (anonymous):

complete ur question...

OpenStudy (anonymous):

what u want to find??

OpenStudy (anonymous):

write the inequality that describes the different color cheryy bombs and roman candles you buy

OpenStudy (anonymous):

ok..

OpenStudy (anonymous):

30x+25y<=100

OpenStudy (anonymous):

if you gound an extra 75.00 in your back pocket, how many more cherry bombs could you buy if you did not need any more roman candles

OpenStudy (anonymous):

two..

OpenStudy (anonymous):

im sorry were you trying to solve that equation for me?

OpenStudy (anonymous):

what??

OpenStudy (anonymous):

the next question is if you found an extra 75.00 in your back pocket how many more cherry bombs could you buy if you did not need any more rome candles

OpenStudy (anonymous):

yes i said two...because 2*30=60<75 but 3*30=90>75

OpenStudy (anonymous):

have u got it??

OpenStudy (anonymous):

yes i think so

OpenStudy (anonymous):

i have another one if you can help

OpenStudy (anonymous):

you take a ball and trhow it up into the air to see how high up it will go. the equation H=-24t (squared) + 135t gives the distance h in feet that the ball is above the ground t seconds after you through it

OpenStudy (anonymous):

how high is the ball after 4 seconds?

OpenStudy (anonymous):

put t=4 here and u get -24*4^2+135*4=156

OpenStudy (anonymous):

is 156 the answer? can you show me the work to get to that?

OpenStudy (anonymous):

put t=4 here and u get -24*4^2+135*4=156

OpenStudy (anonymous):

next question is -- at what height does the ball start to come back down to the ground?

OpenStudy (anonymous):

dipankarstudy are you still ther?

OpenStudy (anonymous):

find H max....using dH/dt =-48t +135=0 t=135/48=45/16 noe put t=45/16 in the equation of the height u will get the required height.

OpenStudy (anonymous):

i dont understand

OpenStudy (anonymous):

do u understand calculus..??

OpenStudy (anonymous):

not really in algebra 2

OpenStudy (anonymous):

next question is how long did it take for the ball to hit the ground?

OpenStudy (anonymous):

find t putting H=0

OpenStudy (anonymous):

t=135/24

OpenStudy (anonymous):

what is the formula for how long did it take the abll to hit the ground?

OpenStudy (anonymous):

dipank?

OpenStudy (anonymous):

are you still ther dipand?

OpenStudy (anonymous):

what is the forula for how long did it take the abll to hit the ground?